# UniTok: A Unified Tokenizer for Visual Generation and Understanding

### v1

@@ Abstract
The representation disparity between visual generation and understanding imposes a critical gap in integrating these capabilities into a single framework. To bridge this gap, we introduce UniTok, a discrete visual tokenizer that encodes fine-grained details for generation while also capturing high-level semantics for understanding. Despite recent studies have shown that these objectives could induce loss conflicts in training, we reveal that the underlying bottleneck stems from limited representational capacity of discrete tokens. We address this by introducing multi-codebook quantization, which divides vector quantization with several independent sub-codebooks to expand the latent feature space, while avoiding training instability caused by overlarge codebooks. Our method significantly raises the upper limit of unified discrete tokenizers to match or even surpass domain-specific continuous tokenizers. For instance, UniTok achieves a remarkable rFID of 0.38 (versus 0.87 for SD-VAE) and a zero-shot accuracy of 78.6% (versus 76.2% for CLIP) on ImageNet. Our code is available at https://github.com/FoundationVision/UniTok.
Visual Tokenizer, Multimodal Generation and Understanding
@@ 1 Introduction
Autoregressive modeling has recently extended its predominance in natural language processing to visual domains. This is characterized by the rapid growth of autoregressive models in image and video generation (Kondratyuk et al., 2023; Tian et al., 2024; Li et al., 2024c), as well as the increasing prevalence of Multimodal Large Language Models (MLLMs) for diverse visual understanding tasks (Liu et al., 2024a; Tong et al., 2024a; Wu et al., 2024c). Given these advancements in individual fields, there is a growing interest in developing unified MLLMs by integrating visual generation and understanding within a single autoregressive framework (Liu et al., 2024b; Team, 2024; Wang et al., 2024).
However, a critical challenge persists in unification: the disparity in visual tokenization for understanding and generation. For instance, the CLIP (Radford et al., 2021) tokenizer, widely used in multimodal understanding, does not naturally fit into generative modeling, which requires precise encoding of fine-grained details. Conversely, the VQVAE (Van Den Oord et al., 2017) tokenizer, which is tailored for autoregressive generation, potentially falls short of capturing high-level semantics crucial for visual understanding (Xie et al., 2024). To address this issue, recent studies attempt to use separate tokenizers for each task (Wu et al., 2024a), yet this increases model complexity without fundamentally bridging the gap in representation. These limitations underscore the need for a unified visual tokenizer to serve both generation and understanding objectives.
Figure 1: (a): The unified tokenizer training paradigm. (b): Comparison with the unified tokenizer VILA-U in terms of ImageNet zero-shot accuracy and reconstruction FID.
Recently, VILA-U (Wu et al., 2024d) has proposed a promising paradigm for unified tokenizers, which integrates CLIP supervision into VQVAE training to complement VQ tokens with text alignment and rich semantics, as illustrated in Figure 1(a). Despite this innovation, it has been observed that the unified tokenizer struggles to converge in training and tends to underperform domain-specific tokenizers. This is commonly attributed to loss conflicts arising from divergent optimization goals (Wu et al., 2024d, a). However, recent studies on visual generation suggest that these objectives might not inherently conflict, as evidenced by the alignment behavior in generative and discriminative representation learning (Yu et al., 2024). This contrast raises an important question: Do reconstruction and contrastive losses truly conflict in unified tokenizer training, or might there be an underlying bottleneck that remains unidentified?
@@
To answer the question, we conduct a comprehensive ablation study on the current unified tokenizer training paradigm (see Figure 3), which yields several intriguing findings. First, we show that removing reconstruction supervision, which equals training a vector-quantized CLIP model, actually does not lead to better understanding performance over unified tokenizers. This observation indicates that the existing gap between unified and CLIP tokenizers mainly arise from vector quantization rather than the conflict of learning objectives. We further establish that this gap stems from both token factorization – which projects tokens into a lower-dimensional space for code index lookup (Yu et al., 2021) – and discretization. These operations are essential for vector quantization but inevitably compromise the expressiveness of visual tokens. We thus identify that the primary bottleneck of unified tokenizers lies in the limited representational capacity of discrete tokens.
To address this limitation, an intuitive approach is to expand the codebook size and latent code dimension of the tokenizer, which enables better approximation of the continuous feature space. However, this in practice could result in low codebook utilization (Zhu et al., 2024) and diminished performance gains (Yu et al., 2023b), while also making the code lookup process more computationally expensive. Drawing inspiration from the divide-and-conquer algorithm, we introduce multi-codebook quantization to alleviate this problem. Specifically, this involves dividing the visual token into multiple chunks and discretizing each with independent sub-codebooks, similar to the multi-head attention mechanism (Vaswani, 2017). Such design effectively scales up the latent code space by increasing the number of sub-codebooks, circumventing the optimization problem associated with large codebooks.
Building upon multi-codebook quantization, we train a unified tokenizer called UniTok to bridge visual generation and understanding. We conduct a thorough evaluation of UniTok’s reconstruction and classification capabilities, and more importantly, its effectiveness in establishing a unified MLLM. Experimental results show that UniTok achieves comparable or even better performance to domain-specific tokenizers, recording an impressive 0.38 reconstruction FID and a 78.6% zero-shot accuracy at $256\times 256$ resolution on ImageNet. The foundational capabilities of UniTok further contribute to strong downstream task performance. Our model sets a new state-of-the-art among unified autoregressive MLLMs on both multimodal understanding and generation benchmarks.
Figure 2: An overview of UniTok. The tokenizer is trained to faithfully reconstruct the input image while aligning its discrete latent features with the text caption. For vector quantization, each visual token is split into multiple chunks, which then undergo code index lookup on corresponding sub-codebooks concurrently.
@@ 2 Related Work
Image Tokenization for Generation. In the domain of visual generation, image tokenization plays an important role in encoding raw pixels into compact latent features for generative modeling (Van Den Oord et al., 2017; Rombach et al., 2022a). Particularly, among a variety of tokenizers, the vector-quantized tokenizer (Van Den Oord et al., 2017) is favored for its discrete latent space and compatibility with autoregressive or masked generative models (Tian et al., 2024; Sun et al., 2024a; Chang et al., 2022; Yu et al., 2023a). The pioneering work VQVAE (Van Den Oord et al., 2017) initially introduced the concept of discretizing continuous tokens by mapping them to the nearest neighbors in a learnable codebook. Built on this, VQGAN (Esser et al., 2021) incorporated perceptual loss (Zhang et al., 2018) and discriminator loss (Karras et al., 2019) to significantly improve the reconstruction quality. Subsequently, ViT-VQGAN (Yu et al., 2021) advanced the framework with the transformer architecture. In recent literature, considerable efforts have been devoted to developing better quantization methods such as residual quantization (Lee et al., 2022) and lookup-free quantization (Yu et al., 2023b), which also constitute a focal point of this paper.
@@
Image Tokenization for Understanding. The unprecedented success of large language models (LLMs) (Wei et al., 2022; Achiam et al., 2023) has catalyzed the development of multimodal large language models (MLLMs) (Liu et al., 2024a; Lin et al., 2024; McKinzie et al., 2024). As a critical component of MLLMs, the selection of an effective vision tokenizer has been the subject of extensive study (Wang et al., 2023; Tong et al., 2024a). A common choice of the vision tokenizer is the pretrained CLIP model (Radford et al., 2021), which undergoes alignment with language during its pretraining phase. While self-supervised learning models, such as DINOv2 (Oquab et al., 2023), are shown to be advantageous at region-level tasks (Ma et al., 2025). Cambrian-1 (Tong et al., 2024a) further demonstrates that MLLMs can benefit from hybrid representations from a mixture of vision encoders. Nonetheless, these tokenizers predominantly encode images into continuous tokens, presenting a challenge for uniformly modeling both vision and text tokens. To address this disparity, several works have explored discretizing CLIP tokens (Ge et al., 2023) or employing VQVAE encoders (Liu et al., 2024b; Xie et al., 2024). However, these approaches have been observed to substantially impair understanding performance of MLLM.
@@
Unified Vision-Language Models. The rise of MLLMs is not limited to the realm of visual understanding. Recent advancements have witnessed an increasing focus on unifying visual generation and understanding within one MLLM (Dong et al., 2023; Wu et al., 2023; Team, 2024; Zhou et al., 2024; Xie et al., 2024; Tong et al., 2024b). Specifically, a line of works employs continuous visual tokenizers for image encoding, and leverages pretrained diffusion models for image synthesis (Dong et al., 2023; Ge et al., 2024; Sun et al., 2024b). This approach inevitably disconnects visual token sampling from the MLLM. In contrast, another stream of research adopts VQVAE models to encode images into discrete tokens (Team, 2024; Wang et al., 2024; Xie et al., 2024; Wu et al., 2024d, b). These tokens are subsequently modeled using the same next token prediction loss that is applied to text tokens, facilitating a unified approach to multimodal learning. However, as reconstruction-oriented VQVAE does not naturally align with the LLM token space, these models typically suffer from degraded visual comprehension capabilities. Our research aligns with the second approach, with a particular focus on the tokenizer design that is suitable for both generation and understanding tasks.
@@ 3 Method
In this section, we introduce UniTok, a unified visual tokenizer well-suited for both generation and understanding tasks. We start with a unified training recipe that integrates reconstruction (VQVAE) and contrastive (CLIP) supervisions (Section 3.1). However, we find that simply combining both training objectives leads to severe performance degradation, which can be mainly attributed to limited representational capacity of discrete tokens (Section 3.2). To this end, we propose multi-codebook quantization and attention factorization to enhance the latent feature space and derive unified visual representations (Section 3.3). An overview of the framework is presented in Figure 2.
@@ 3.1 Unified Supervision
Visual generation and understanding typically impose distinct demands on the visual tokenizer. For instance, generation emphasizes lossless compression for accurate reconstruction, whereas understanding prioritizes semantically meaningful and discriminative features. To accommodate both requirements, we jointly train the tokenizer with (i) a VQVAE-based reconstruction loss to preserve low-level information, and (ii) an image-text contrastive loss that enhances high-level semantics of the features.
To be specific, the VQVAE-based loss term $\mathcal{L}_{\text{recon}}$ consists of a pixel-level reconstruction loss $\mathcal{L}_{\text{R}}$, a perceptual loss $\mathcal{L}_{\text{P}}$ based on the LPIPS metric (Zhang et al., 2018), a discriminator loss $\mathcal{L}_{\text{G}}$ to enhance reconstruction fidelity (Karras et al., 2019), and a vector quantization loss $\mathcal{L}_{\text{VQ}}$ to minimize distance between the encoder output and its nearest code entry. It is denoted as:
@@
where $\lambda$ is the weight factor for the corresponding loss term. The image-text contrastive loss term $\mathcal{L}_{\text{contra}}$ is basically the same as in CLIP (Radford et al., 2021). Therefore, the final loss term of UniTok can be written as:
@@ 3.2 Quantization Bottleneck
Despite being augmented with CLIP supervision, we find that the unified tokenizer still exhibits unsatisfactory performance in visual understanding tasks, significantly lagging behind the commonly used CLIP tokenizer. To figure out the underlying causes of this underperformance, we break down the key components involved in training a unified tokenizer, as illustrated in Figure 3. Starting with the CLIP baseline, we provide a step-by-step walk-through and ablation of all changes in the following paragraphs.
Figure 3: Roadmap to build UniTok. The blue bars illustrate the progressive changes in VQA performance from the CLIP tokenizer to the unified tokenizer, while the purple bars represent the proposed improvements in UniTok. The VQA score is measured using the average accuracy across the VQAv2, GQA, TextVQA, and POPE benchmarks. All models are trained from scratch on 512m image-text pairs from DataComp.
Factorization. Modern VQ-tokenizers typically project continuous tokens to a lower-dimensional latent space for code index lookup (e.g. from 768-d to 16-d), known as token factorization (Yu et al., 2021). This increases the relative density of codes by compressing the latent code space, thereby reducing quantization error. To evaluate the impact of factorization in CLIP training, we add two linear projection layers on top of the CLIP vision encoder (right before average pooling), which transforms tokens from 768-d to 16-d and then back to 768-d. Notably, vector quantization and reconstruction supervision are not included at this stage. Surprisingly, it turns out that this channel compression operation significantly compromises the expressiveness of tokens, leading to severe performance degradation in downstream VQA tasks.
@@
Reconstruction Supervision. Finally, we integrate reconstruction losses into the training process to build a unified tokenizer, as outlined in Section 3.1. Previous literature suggests that loss conflict between VQVAE and CLIP is a major cause of performance degradation in joint training (Wu et al., 2024d). We observe a similar phenomenon where joint training results in sub-optimal ImageNet zero-shot classification accuracy and reconstruction FID compared to specialized training. However, surprisingly, we find that this degradation has negligible impacts on downstream understanding performance. Moreover, the degradation in classification accuracy and reconstruction FID diminishes after we improve the quantization methods (detailed in the next section). Based on these observations, we speculate that the perceived loss conflict is only a superficial issue, and the primary cause of the underperformance lies in the limited representational capacity of discrete tokens.
@@ 3.3 UniTok
A straightforward solution to breaking the quantization bottleneck could be increasing the codebook size and the latent code dimension. However, current studies on VQVAE tokenizers suggest that there is diminishing gain in scaling and the performance saturates after the codebook size reaches 16k (Yu et al., 2023b; Sun et al., 2024a). Continuing expansion results in a substantial portion of codes being rarely used or becoming ‘dead’ during training, which negatively impacts downstream task performance (Yu et al., 2021). To address this issue, we propose multi-codebook quantization and attention factorization in the following paragraphs.
@@
Attention factorization. Existing VQ methods usually employ linear or convolutional projection layers for token factorization. But as shown in Section 3.2, this over-simplified design fails to preserve rich semantics in original tokens, leading to degraded understanding performance. To alleviate this problem, we suggest adapting the multi-head attention modules for factorization, as illustrated in Figure 4. Despite its simplicity, we find this design effectively strengthens the representational power of factorized tokens. Notably, to ensure compatibility with autoregressive generation, the factorization blocks are configured with causal attention.
Figure 4: Modified attention blocks for factorization. Modules in yellow indicate a change in the number of channels.

### v2

@@ Abstract
Visual generative and understanding models typically rely on distinct tokenizers to process images, presenting a key challenge for unifying them within a single framework. Recent studies attempt to address this by connecting the training of VQVAE (for autoregressive generation) and CLIP (for understanding) to build a unified tokenizer. However, directly combining these training objectives has been observed to cause severe loss conflicts. In this paper, we show that reconstruction and semantic supervision do not inherently conflict. Instead, the underlying bottleneck stems from limited representational capacity of discrete token space. Building on these insights, we introduce UniTok, a unified tokenizer featuring a novel multi-codebook quantization mechanism that effectively scales up the vocabulary size and bottleneck dimension. In terms of final performance, UniTok sets a new record of 0.38 rFID and 78.6% zero-shot accuracy on ImageNet. Besides, UniTok can be seamlessly integrated into MLLMs to unlock native visual generation capability, without compromising the understanding performance. Additionally, we show that UniTok favors cfg-free generation, reducing gFID from 14.6 to 2.5 on ImageNet 256$\times$256 benchmark. GitHub: https://github.com/FoundationVision/UniTok.
Figure 1: The major challenge in unified tokenizer training. CLIP supervision cannot be easily incorporated into VQVAE training – This provides only marginal improvements in understanding performance, while drastically degrading reconstruction FID.
@@ 1 Introduction
The advent of GPT-4o [34] highlights the immense potential of Multimodal Large Language Models (MLLMs) with native visual generation capabilities [12, 47, 63, 74, 59]. These unified models offer precise control in multimodal interactions, enabling exceptional fluency in tasks such as multi-turn image editing and visual in-context learning. However, a fundamental dilemma remains in the choice of visual tokenizers for unified MLLMs – e.g., the CLIP [39, 72] tokenizer excels in multimodal understanding but complicates generative modeling due to its high-dimensional, continuous feature space; Conversely, the discrete VQVAE [8] tokenizer fits autoregressive generation but struggles to capture essential semantics for understanding [63].
In this work, we aim to design a unified visual tokenizer to bridge the gap in multimodal generation and understanding. Intuitively, this can be achieved by integrating CLIP supervision into VQVAE training, resulting in a discrete tokenizer capturing both fine-grained details and high-level semantics. However, we empirically find this training recipe confronts severe convergence issues [62] and largely falls behind the CLIP baseline in multimodal understanding (Figure 1). While prior studies commonly attribute these challenges to conflicts between semantic and pixel-level feature learning [62, 38, 59], recent progress in visual generation suggests the opposite, showing semantic regularization could benefit tokenizers in reconstruction-oriented training [64, 4, 19]. Such disparity motivates us question: Do reconstruction and semantic losses truly conflict in tokenizer training?
@@
To study the problem, we conduct a comprehensive ablation on the unified tokenizer training paradigm (Figure 3), which yields several intriguing findings: First, we show that removing reconstruction supervision, which leads to a vector-quantized CLIP model, does not improve understanding performance compared to the unified tokenizer. This observation indicates that the performance gap between unified and CLIP tokenizers mainly arises from vector quantization, rather than conflicts between learning objectives; Further analysis reveals that this gap is driven by two key factors: token factorization, which projects tokens into a lower-dimensional space for code index lookup [66], and discretization. These operations are essential for vector quantization but inevitably compromise the expressiveness of visual tokens. We thus argue that the primary bottleneck of unified tokenizers lies in the limited representational capacity of discrete token space.
In light of the issue, we consider expanding the vocabulary size and latent code dimension, which allows for a closer approximation of the continuous feature space. However, extensive studies have shown that doing so could result in low codebook utilization [75, 66] and diminishing performance gains [68]. To address this, we introduce multi-codebook quantization to partition the visual token into several chunks, each discretized using a small, separate sub-codebook, akin to the multi-head attention mechanism [55]. This design exponentially scales the vocabulary size with the number of sub-codebooks, while avoiding the optimization problems of large monolithic codebooks. Besides, we replace traditional linear projection layers with adapted attention modules for token factorization, which is observed to consistently improve training stability and understanding performance.
Building upon these techniques, we train a unified tokenizer called UniTok to bridge visual generation and understanding. Through extensive experiments, we demonstrate that UniTok achieves comparable or even better performance to domain-specific tokenizers: On ImageNet evaluation, UniTok records an impressive 0.38 reconstruction FID and 78.6% zero-shot accuracy at 256$\times$256 resolution; In building unified MLLMs, UniTok enables the MLLM with native visual generation capabilities while maintaining decent understanding performance. It outperforms the Liquid [60] baseline with a VQGAN tokenizer by 5.5% on VQAv2 [13], 9.2% on TextVQA [43], and 339 points on MME [65]; In addition, we demonstrate that semantic supervision leads to improved latent space structure for autoregressive generation, i.e., for class-conditional image generation on ImageNet 256$\times$256, UniTok significantly reduces generation FID without classifier-free guidance from 14.6 to 2.5 under the LlamaGen [44] framework, which aligns with recent findings in diffusion modeling [19, 4, 64].
@@ 2 Related Work
In the domain of visual generation, image tokenization plays an important role in encoding raw pixels into compact latent features for generative modeling [54, 40]. Among a variety of tokenizers, the vector-quantized tokenizer [54] is favored for its discrete latent space and compatibility with autoregressive or masked generative models [49, 44, 3, 67]. The pioneering work VQVAE [54] initially introduced the concept of discretizing continuous tokens by mapping them to the nearest neighbors in a learnable codebook. Built on this, VQGAN [8] added perceptual loss [73] and discriminator loss [16] to improve the reconstruction quality. ViT-VQGAN [66] subsequently advanced the framework with the transformer architecture. In recent literature, considerable efforts have been devoted to developing better quantization methods such as residual quantization [18] and lookup-free quantization [68], which also constitute a focal point of this paper.
Image Tokenization for Understanding.
@@
The unprecedented success of large language models (LLMs) [58, 1, 52, 48] has catalyzed the development of multimodal large language models (MLLMs) [28, 25, 32]. As a critical component of MLLMs, the selection of an effective vision tokenizer has been the subject of extensive study [56, 50]. A common choice of the vision tokenizer is the pretrained CLIP model [39], which undergoes alignment with language during its pretraining phase. While self-supervised learning models, such as DINOv2 [35], are shown to be advantageous at region-level tasks [30]. However, these tokenizers predominantly encode images into a continuous feature space, presenting challenges for uniformly modeling both vision and text tokens. To address this, some works have explored discretizing CLIP tokens [10] or employing VQVAE encoders [27, 63]. Yet, these methods have been observed to substantially impair understanding performance of MLLMs.
Unified Vision-Language Models.
@@
The rise of MLLMs is not limited to the realm of visual understanding. Recent advancements have witnessed an increasing focus on unifying visual generation and understanding within one MLLM [7, 61, 47, 74, 63, 51, 21]. Specifically, a line of works employs continuous visual tokenizers for image encoding, and leverages pretrained diffusion models for image synthesis [7, 11, 45]. This approach inevitably increases model complexity and disconnects the visual sampling process from the MLLM. In contrast, another stream of research adopts VQVAE models to encode images into discrete tokens [47, 57, 63, 62, 60]. These tokens are subsequently modeled using the same cross-entropy loss that is applied to text tokens, facilitating a unified approach to multimodal learning. However, as reconstruction-oriented VQVAE does not naturally align with the LLM token space, these models typically suffer from degraded visual comprehension capabilities. Our research aligns with the second approach, with a particular focus on the tokenizer design that is suitable for both generation and understanding tasks.
@@ 3 Method
In this section, we introduce UniTok, a unified tokenizer well-suited for both visual generation and understanding tasks. We start with a unified training recipe that integrates reconstruction (VQVAE) and semantic (CLIP) supervisions (Section 3.1). However, we find that simply combining both training objectives leads to severe performance degradation, which can be mainly attributed to limited representational capacity of discrete tokens (Section 3.2). To this end, we propose multi-codebook quantization and attention projection to enhance the latent feature space and derive unified visual representations (Section 3.3). An overview of the framework is presented in Figure 2.
Figure 2: An overview of UniTok. The tokenizer is trained to reconstruct the input image while aligning its discrete latent features with the text caption. For vector quantization, each visual token is split into multiple chunks, which then undergo code index lookup on corresponding sub-codebooks.
@@ 3.1 Unified Supervision
Visual generative and understanding models typically impose distinct demands on the visual tokenizers. For instance, generation emphasizes precise encoding of the visual signals, whereas understanding prioritizes capturing high-level semantics. To accommodate both requirements, we jointly train the tokenizer with (i) a VQVAE-based reconstruction loss to preserve low-level information, and (ii) an image-text contrastive loss that enhances high-level semantics of the features.
To be specific, the VQVAE-based loss term $\mathcal{L}_{\text{recon}}$ consists of a pixel-level reconstruction loss $\mathcal{L}_{\text{R}}$, a perceptual loss $\mathcal{L}_{\text{P}}$ based on the LPIPS metric [73], a discriminator loss $\mathcal{L}_{\text{G}}$ to enhance reconstruction fidelity [16], and a vector quantization loss $\mathcal{L}_{\text{VQ}}$ to minimize distance between the encoder output and its nearest code entry. It is denoted as:
@@
where $\lambda$ is the weight factor for the corresponding loss term. The image-text contrastive loss term $\mathcal{L}_{\text{contra}}$ is basically the same as in CLIP [39]. Therefore, the final loss term can be written as:
@@ 3.2 Quantization Bottleneck
Despite being augmented with CLIP supervision, we find that the unified tokenizer exhibits unsatisfactory performance in visual understanding tasks, significantly lagging behind the commonly used CLIP tokenizer. To figure out the underlying cause of this underperformance, we break down the key components involved in training a unified tokenizer, as illustrated in Figure 3. Starting with the CLIP baseline, we provide a step-by-step walk-through of all changes in following paragraphs.
Figure 3: Roadmap from CLIP to UniTok. It is observed that major degradation in understanding performance comes from token factorization and discretization, rather than reconstruction supervision. The proposed multi-codebook quantization and attention projection effectively address this by scaling up the vocabulary size and bottleneck dimension. The VQA score is measured using the average score across the VQAv2, GQA, TextVQA, and POPE benchmarks. All tokenizers are trained from scratch on 512m image-text pairs from DataComp.
Factorization. Modern VQ-tokenizers typically project continuous tokens to a lower-dimensional latent space for code index lookup (e.g. from 768-d to 8-d), known as token factorization [66]. This increases the relative density of codes by compressing the latent code space, thereby reducing quantization error. To evaluate the impact of factorization in CLIP training, we add two linear projection layers on top of the CLIP vision encoder (right before average pooling), which transforms tokens from 768-d to 16-d and then back to 768-d. Notably, vector quantization and reconstruction supervision are not included at this stage. Surprisingly, it turns out that this channel compression operation significantly compromises the expressiveness of tokens, leading to severe performance degradation in downstream VQA tasks.
@@
Reconstruction Supervision. Finally, we integrate reconstruction losses into the training process to build a unified tokenizer, as outlined in Section 3.1. Previous literature suggests that loss conflict between VQVAE and CLIP is a major cause of performance degradation in joint training [62]. We observe a similar phenomenon where joint training results in sub-optimal ImageNet zero-shot classification accuracy and reconstruction FID compared to specialized training. However, surprisingly, we find that this degradation has negligible impacts on downstream understanding performance. Moreover, the degradation in classification accuracy and reconstruction FID diminishes after we improve the quantization methods (detailed in the next section). Based on these observations, we speculate that the perceived loss conflict is only a superficial issue, and the primary cause of the underperformance lies in the limited representational capacity of discrete tokens.
@@ 3.3 UniTok
A straightforward solution to breaking the quantization bottleneck could be increasing the codebook size and the latent code dimension. However, current studies on VQVAE tokenizers suggest that there is diminishing gain in scaling and the performance saturates after the codebook size reaches 16k [68, 44]. Continuing expansion results in a substantial portion of codes being rarely used or becoming ‘dead’ during training, which negatively impacts downstream task performance [66]. To address this, we propose multi-codebook quantization and attention projection in the following paragraphs.
@@
Discussions. MCQ shares a similar concept with residual quantization (RQ) [18] in using multiple codes to quantize a token, but differs fundamentally in design philosophy: RQ follows a coarse-to-fine quantization order, whereas MCQ adopts a divide-and-conquer strategy. This distinction gives MCQ unique advantages when operating in high-dimensional latent spaces, where codes tend to become increasingly sparse. For instance, with a latent dimension of 64-d, we observe that MCQ’s quantization loss is 15 to 45 times lower than that of RQ. This is because MCQ partitions the original latent space into multiple low-dimensional subspaces for quantization. Our ablation study in Table 8 further confirms the superiority of MCQ in unified tokenizer training.
Attention projection. Existing VQ methods usually employ linear or convolutional projection layers for token factorization. But as shown in Figure 3, this over-simplified design fails to preserve rich semantics when compressing the feature dimensions, leading to degraded understanding performance. To alleviate this problem, we suggest adapting the multi-head attention modules for factorization. Specifically, instead of concatenating features from multiple heads after the attention calculation, we replace the concatenation operation with average pooling to realize channel compression. Figure 5 provides a detailed illustration of the adaptation. Despite its simplicity, we find this design effectively strengthens the representational power of factorized tokens and stabilizes training.